# Unifying Large Language Models and Knowledge Graphs: A Roadmap

### v1

@@ 7.4 Multi-Modal LLMs for KGs
Current knowledge graphs typically rely on textual and graph structure to handle KG-related applications. However, real-world knowledge graphs are often constructed by data from diverse modalities [255, 256, 101]. Therefore, effectively leveraging representations from multiple modalities would be a significant challenge for future research in KGs. One potential solution is to develop methods that can accurately encode and align entities across different modalities. Recently, with the development of multi-modal LLMs [257, 100], leveraging LLMs for modality alignment holds promise in this regard. But, bridging the gap between multi-modal LLMs and KG structure remains a crucial challenge in this field, demanding further investigation and advancements.
@@ 8 Conclusion
Unifying large language models (LLMs) and knowledge graphs (KGs) is an active research direction that has attracted increasing attention from both academia and industry. In this article, we provide a thorough overview of the recent research in this field. We first introduce different manners that integrate KGs to enhance LLMs. Then, we introduce existing methods that apply LLMs for KGs and establish taxonomy based on varieties of KG tasks. Finally, we discuss the challenges and future directions in this field. We hope this article can provide a comprehensive understanding of this field and advance future research.

### v3

@@ 7.4 Multi-Modal LLMs for KGs
Current knowledge graphs typically rely on textual and graph structure to handle KG-related applications. However, real-world knowledge graphs are often constructed by data from diverse modalities [257, 258, 99]. Therefore, effectively leveraging representations from multiple modalities would be a significant challenge for future research in KGs [259]. One potential solution is to develop methods that can accurately encode and align entities across different modalities. Recently, with the development of multi-modal LLMs [260, 98], leveraging LLMs for modality alignment holds promise in this regard. But, bridging the gap between multi-modal LLMs and KG structure remains a crucial challenge in this field, demanding further investigation and advancements.
@@ 8 Conclusion
Unifying large language models (LLMs) and knowledge graphs (KGs) is an active research direction that has attracted increasing attention from both academia and industry. In this article, we provide a thorough overview of the recent research in this field. We first introduce different manners that integrate KGs to enhance LLMs. Then, we introduce existing methods that apply LLMs for KGs and establish taxonomy based on varieties of KG tasks. Finally, we discuss the challenges and future directions in this field. We envision that there will be multiple stages (milestones) in the roadmap of unifying KGs and LLMs, as shown in Fig. 26. In particular, we will anticipate increasing research on three stages: Stage 1: KG-enhanced LLMs, LLM-augmented KGs, Stage 2: Synergized LLMs + KGs, and Stage 3: Graph Structure Understanding, Multi-modality, Knowledge Updating. We hope that this article will provide a guideline to advance future research.
Acknowledgments
This research was supported by the Australian Research Council (ARC) under grants FT210100097 and DP240101547 and the National Natural Science Foundation of China (NSFC) under grant 62120106008.
Appendix A Pros and Cons for LLMs and KGs
In this section, we introduce the pros and cons of LLMs and KGs in detail. We summarize the pros and cons of LLMs and KGs in Fig. 1, respectively.
LLM pros.
  • General Knowledge [11]: LLMs pre-trained on large-scale corpora, which contain a large amount of general knowledge, such as commonsense knowledge [264] and factual knowledge [14]. Such knowledge can be distilled from LLMs and used for downstream tasks [265].
  • Language Processing [12]: LLMs have shown great performance in understanding natural language [266]. Therefore, LLMs can be used in many natural language processing tasks, such as question answering [4], machine translation [5], and text generation [6].
  • Generalizability [13]: LLMs enable great generalizability, which can be applied to various downstream tasks [267]. By providing few-shot examples [59] or finetuning on multi-task data [3], LLMs achieve great performance on many tasks.
LLM cons.
  • Implicit Knowledge [14]: LLMs represent knowledge implicitly in their parameters. It is difficult to interpret or validate the knowledge obtained by LLMs.
  • Hallucination [15]: LLMs often experience hallucinations by generating content that while seemingly plausible but are factually incorrect [268]. This problem greatly reduces the trustworthiness of LLMs in real-world scenarios.
  • Indecisiveness [16]: LLMs perform reasoning by generating from a probability model, which is an indecisive process. The generated results are sampled from the probability distribution, which is difficult to control.
  • Black-box [17]: LLMs are criticized for their lack of interpretability. It is unclear to know the specific patterns and functions LLMs use to arrive at predictions or decisions.
  • Lacking Domain-specific/New Knowledge [18]: LLMs trained on general corpus might not be able to generalize well to specific domains or new knowledge due to the lack of domain-specific knowledge or new training data.
KG pros.
  • Structural Knowledge [19]: KGs store facts in a structural format (i.e., triples), which can be understandable by both humans and machines.
  • Accuracy [20]: Facts in KGs are usually manually curated or validated by experts, which are more accurate and dependable than those in LLMs.
  • Decisiveness [21]: The factual knowledge in KGs is stored in a decisive manner. The reasoning algorithm in KGs is also deterministic, which can provide decisive results.
  • Interpretability [22]: KGs are renowned for their symbolic reasoning ability, which provides an interpretable reasoning process that can be understood by humans.
  • Domain-specific Knowledge [23]: Many domains can construct their KGs by experts to provide precise and dependable domain-specific knowledge.
  • Evolving Knowledge [24]: The facts in KGs are continuously evolving. The KGs can be updated with new facts by inserting new triples and deleting outdated ones.
KG cons.
  • Incompleteness [25]: KGs are hard to construct and often incomplete, which limits the ability of KGs to provide comprehensive knowledge.
  • Lacking Language Understanding [33]: Most studies on KGs model the structure of knowledge, but ignore the textual information in KGs. The textual information in KGs is often ignored in KG-related tasks, such as KG completion [26] and KGQA [43].
  • Unseen Facts [27]: KGs are dynamically changing, which makes it difficult to model unseen entities and represent new facts.